# Atacama Large Aperture Submillimeter Telescope (AtLAST) Science: Solar and stellar observations

### v2

@@ Introduction
Measuring the solar radius at subterahertz frequencies (submm/mm) allows one to probe the solar atmosphere, since these measurements show the height above the photosphere at which most of the emission at a determined observation frequency is generated (Menezes & Valio 2017). Changes in the solar radius show it can be modulated with the 11-year activity cycle (mid-term variations) as well as longer periods (long-term variations). The equatorial radius time series was found to be positively correlated to the solar cycle, since the equatorial regions are more affected by the increase of active regions during solar maxima, making the solar atmosphere warmer in these regions; on the other hand, the anticorrelation between polar radius time series and the solar activity proxies could be explained by a possible increase of polar limb brightening during solar minima (Menezes et al. 2021). In the context of the study on the subterahertz solar atmosphere, spicules can affect the measurement of the solar radius and cause the solar limb to appear more diffuse, which can lead to an overestimation of the solar radius. In Menezes et al. (2022), the discrepancies between measured limb brightening values and model predictions highlight the need for further studies to improve our understanding of the solar atmosphere. Therefore, it is important to take into account the presence of spicules when measuring the solar radius and limb brightening at different frequencies. Currently, however, there are no full-disk observations of the Sun in the mm range apart from occasional ALMA TP maps a few times a year.
@@
Please note that the Sun’s rotation depends on latitude with a period of down to 24 days at the equator (see, e.g., Schou et al. 1998; Snodgrass 1983). Any long-lived feature like an Active Region would thus move across the visible disk of the Sun in less than two weeks. However, at the resolution anticipated for AtLAST, features like Active Regions and the sunspots (Solanki 2003) therein would evolve significantly on a daily timescale. This use case would therefore be addressed best via full-disk mapping of the Sun at multiple frequency bands, once per day. Already an initial 1-month campaign covering a full solar rotation would allow for addressing the fundamental unsolved problems mentioned above including the nature of coronal and chromospheric heating problems and solar flares, whereas much longer sequences are required for addressing the solar activity cycle.
@@
New insights regarding the Sun’s activity variations and thus the long-term evolution of our host star could be transferred to other stars. This Sun-as-a-star approach has recently received increasing attention in the context of next-generation exoplanet observations for which it is crucial to separate observable exoplanet signatures from the host star’s “background radiation”.
@@
The Sun, which can be observed spatially resolved, serves as a fundamental reference case for studying the dynamic nature of stellar atmospheres and solar/stellar activity including the occurrence of flares (see Sect. 2.2). In addition to serendipitous detection as a by-product of other observations and surveys, dedicated stellar observations are essential for understanding the structure and activity of stellar atmospheres in general. The sample by Mohan et al. (2021) demonstrates the potential of stellar mm observations and, at the same time, illustrates the scarcity of suitable observations. AtLAST could make substantial contributions in this regard, either through targeted observations or also through coincidental (or even archival) detections. In particular, observations of stellar flares in the (sub-)mm range are extremely rare (MacGregor et al. 2018, 2020), and yet are crucial for unveiling the physics of these events including particle acceleration and plasma heating and for evaluating the impact of flares on the habitability of exoplanets (see, e.g., Lammer et al. 2003; Segura et al. 2010; Konings et al. 2022; Bellotti et al. 2023, and references therein). At millimeter wavelengths, the flux density of the few detected stellar flares tends to be 10 - 1000 times higher than the quiescent flux. Assuming a typical quiet chromospheric temperature of $\sim 10^{5}$ K (at 100 GHz) as seen in cool stars (Mohan et al. 2021), the flux density during a flare can vary in the range $\sim$ 0.5 - 50 Jy for a star at a distance of 1 pc. It is important to emphasise that the flux varies on timescales of a few seconds to minutes. Observing the temporal evolution of a flare, which is essential for exploring the physical mechanisms, thus requires time sequences with accordingly high temporal cadence. Such data sets are typically not generated from a serendipitous detection or as part of a survey but require a dedicated sit-and-stare observation of the same region in the sky.
@@
The full potential of flare observations at millimeter wavelengths is only unlocked through coordinated, strictly co-simultaneous observations with other telescopes covering as large a wavelength range across the electromagnetic spectrum as possible. Such coordinated campaigns, which have become common in multi-messenger astronomy and are basically standard procedure for observations of the Sun, demand for PI-driven operations. The scarcity of such coordinated stellar observing campaigns is among the main reasons why there is still no confident detection of a stellar coronal mass ejection (CME) – an essential component of exo-space weather. The success of such campaigns would be greatly enhanced by AtLAST’s large FOV and high angular resolution, thus being able to observe multiple stellar targets at the same time while separating the flare source(s) from other components such as close-by active companion stars or disks, making AtLAST ideal for time-resolved coordinated observations of stellar flares. See also the time-domain science cases for AtLAST (Orlowski-Scherer & et al. 2024).
@@
Figure 3: The ALMA sunspot mosaic (left) that was taken during the ALMA SV campaign in on December 18, 2015 in Band 6 230 GHz at original interferometric resolution (left) and after convolution with the expected AtLAST beam (right). For reference, the ALMA single-dish beam (dashed line) and the AtLAST beam (solid line) are plotted in the lower left corner.
@@
There are two major constraints for solar observations at millimeter wavelengths: (i) As the Sun is by far the brightest millimeter source in the sky, sensitivity is not an issue and very short integration times and thus even on-the-fly scanning is possible. (ii) The Sun evolves on short time scales which prohibits long integration/scan times. Imaging of solar data thus needs to be done as snapshot imaging or over short time windows that should not exceed 1 min for the expected resolution (on the Sun, temporal and spatial scales tend to correlate). As will be detailed in this section, the combination of high angular resolution, high temporal cadence (or at least short integration time), wide frequency coverage, and large dynamic range are crucial for meaningful solar observations with AtLAST (Mroczkowski et al. 2024). In addition, full-Stokes polarimetry would greatly enhance AtLAST diagnostic potential for the Sun.
@@
High spatial resolution is of utmost importance for solar observations as much of the relevant dynamics occurs on small spatial scales. Achieving sufficiently high resolution is challenging at millimeter wavelengths, in particular for single-dish telescopes such as AtLAST. While AtLAST’s aperture sets a limit, the achievable resolution is still in the relevant range and superior to any previous single-dish observation of the Sun. With an aperture of 50 m, the achievable angular resolution would be 6.5” at a wavelength of 1.3 mm (ALMA Band 6) and down to just 1.5” at 0.3 mm (ALMA Band 10) (see Fig.  4).
@@
techniques (with a single beam) are used for creating maps of the whole solar disk or limited regions by, e.g, ALMA (White et al. 2017), and could be implemented for AtLAST, too (Kirkaune et al. in prep.). The scanning motion could be accomplished by moving the primary or secondary reflector. Scanning the whole disk with an ALMA TP antenna in a double-circle pattern (Phillips et al. 2015) takes on the order of 10 min. See Fig. 1 for examples. The roughly 4 times larger aperture of AtLAST compared to an ALMA 12-m antenna results in an accordingly much smaller ($\sim$6 %) beam area. Scanning the whole disk of the Sun would take considerably longer, which would be sufficient for synoptic studies with just one scan per day but prohibitive for science cases that rely on high temporal cadence. In the latter case, scans would then need to be limited to smaller region, although a compromise between size of the mapped region and overall cadence can be made. In either case, for a time series, the scanning could be continuous.
@@
would significantly enhance the diagnostic possibilities for solar observations with AtLAST. Even for a modest setup with a detector with 1000 pixels in a concentric setup with no overlap, spaced by one beam width, the resulting instantaneously covered region on the Sun (hereafter referred to as the instantaneous field of view, FOV) would cover a region with a radius of about 36 beams. While a single AtLAST beam would be a factor $50/12$ smaller than the beam of a 12 m ALMA antenna, using 1000 pixels would effectively make the instantaneous FOV $36\times\frac{12}{50}\approx 9$ times larger. As shown in Fig. 4, the diameter of the FOV would range from $\sim 45$” at 950 GHz (ALMA Band 10) to $\sim 450$” at 100 GHz (ALMA Band 3). Such a FOV would cover a complete sunspot even at the highest frequencies and whole Active Regions at the lower frequencies. The angular resolution of the final imaging products could be increased by means of a rapid small scan pattern (on scales comparable to the beam size and below) that compensates for the above described pixel spacing. The diameter of the instantaneous FOV would be halved again for a denser pixel setup that allows for Nyquist sampling, which then would not require the mentioned scan pattern. Please see Sect. 3.3 for a detailed discussion of the detector requirements for solar observations.
@@
might offer means to further increase AtLAST angular resolution. However, as the formation height range of the mm continuum depends on the frequency, only a moderate frequency range should be considered. Depending on the receiver design, multi-frequency synthesis could be an option that could be applied flexibly on AtLAST data sets.
@@
The Sun is by far the brightest millimeter source in the sky with typical brightness temperatures of several thousand Kelvin. Consequently, sensitivity is not a concern and the integration time can be very low, allowing for ultra-high cadence, which is of essence for a range of science cases. The integration time can be much below 1 s depending on possible attenuation strategies (e.g., solar filter or detuning). Important for solar observations, however, is a large dynamic range. ALMA observations in Band 3 and 6 show a wide range of brightness temperatures from as low as 3000 K to above 13 000 K (see, e.g., data sets in the Solar ALMA Science Archive, Henriques et al. 2022). In general, mean brightness temperature of the Sun depends on the frequency because the atmospheric height range of the mapped layer depends on frequency and thus on the thermal stratification of the solar atmosphere. In principle, the highest ALMA receiver bands map the transition between the upper photosphere and lower chromosphere (i.e., the classical “temperature minimum” region) with comparatively low temperatures, whereas the lowest ALMA bands map radiation emerging from the on average much hotter upper chromosphere close to the transition region with its steep jump in temperature. However, as mentioned in Sect. 2.2, solar flares are a key science case for solar observations with AtLAST. Depending on the strength (i.e. overall energy release) of a flare, plasma temperatures exceed $10^{6}$ K or even $10^{7}$ K. A flare thus emits orders of magnitude more radiation than the average (quiescent) Sun as discussed above. It should be noted that the corresponding peak brightness temperature also depends on the observing frequency and the beam size, which is typically too large to resolve the source. From observations of a (very strong) X-class flare with the 1.5m-aperture Solar Submillimeter Telescope (SST, Kaufmann et al. 2008) at 212 GHz an excess brightness temperature of $1.75\times 10^{6}$ K for an assumed source size of 25”, which would remain unresolved (Giménez de Castro et al. 2018). In comparison, AtLAST would have an angular resolution of $\sim$7” at 212 GHz and would thus be able to resolve the source and thus brightness temperatures in excess of $10^{6}$ K. At shorter wavelengths in the sub-mm range, we rely on simulations to estimate the expected brightness of flares (see Figure 2).
@@
Magnetic fields play an essential role in the Sun’s structure, dynamics, and activity. Their impact is clearly visible in Active Regions most notably in the form of sunspots and as large-scale prominences but also the Sun’s Quiet Regions harbour magnetic fields in various configurations and strengths. The latter becomes immediately obvious from high-resolution H observations (see, e.g., Rutten et al. 2008; Rouppe van der Voort et al. 2023). High-resolution magnetograms obtained from the spectral lines in the photosphere (i.e., the Sun’s surface layer) have become a staple in contemporary solar physics relevant to a large range of scientific topics but often also just for context. Reliably measuring the magnetic field in the solar chromosphere above, i.e. the layer that would be observed with AtLAST, is still a current technological challenge, again with much diagnostic potential. The measured polarization state provides a measure for the longitudinal component of the magnetic field vector in the same layer as the continuum radiation. In the same way, a scan through wavelength (and thus through formation height) can be used to reconstruct the three-dimensional magnetic field structure in the solar chromosphere. Measuring the magnetic field in this layer is in itself a hot topic with potentially ground-breaking results.
Early observations with ALMA at a frequency of 100 GHz, which were obtained as science verification data in Band 3 (ADS/JAO.ALMA#2011.0.00011.E), indicate that the circular polarization degree is at least 1% . This value is considered a minimum, as the observations were carried out with the 12m-Array only (Shimojo et al. 2024). First regular ALMA observations of the Sun with full polarisation in Band 3 are scheduled for Cycle 10. Important lessons are expected from these observations, which would have implications for full polarisation observations with AtLAST. In general, circular polarization levels of up to 5% are expected for the Sun at mm wavelengths.
@@
In order to address the aforementioned key science cases (see Sect. 2), it would be best to combine 1) long-term synoptic observations with a very low daily load on observing time with 2) campaign-based observations. The synoptic observations could consist of one daily full-disk map or preferably a 10 min long time sequence of maps across a wide predefined frequency range. The resulting long-term data set would be of essential value for studying the solar cycle and all related variations imprinted in the Sun’s chromosphere with fundamental implications for stellar activity cycles and their impact on the habitability of exoplanets (White et al. 2023). The campaign-based observations would be more adjusted towards the requirements of individual science cases. The scheduling for observations that require the existence of an Active Region could be planned with just a few days ahead as the co-ordination with other space-borne and ground-based observatories would boost the scientific impact of such AtLAST data. Alternatively, a ToO mode (also possibly confined to a campaign) would ensure to observe flares whenever a suitable Active Region appears.
@@
A field-of-view (FOV) that would cover the whole Sun plus a margin around, which is important to include the Sun’s outer atmosphere, has a diameter of $\sim 2000$”. Scanning such a region with the comparatively small beam size of AtLAST would require long scan times. Therefore we expect that multi-pixel detectors will be needed for solar disk mapping as well as other non-solar large FOV studies. The corresponding time needed for scanning the whole FOV would be reduced basically by the number of simultaneously used beams (pixels) and thus the resulting size of the instantaneously covered FOV (see also Sect. 3.1.1). The required scan time for the whole solar disk would scale accordingly with respect to the single-dish Total Power scans with ALMA antennas, which take about $\sim 10$ min. Consequently, a full-disk scan at the lower frequencies could potentially be completed in a few minutes or less, aided by the anticipated high slew speed of 3 deg/s. A smaller number of pixels, e.g. $\sim 250$ pixels, would directly affect the covered FOV, resulting in correspondingly longer scan times. Please note that very short integration times are sufficient as the Sun is the brightest mm source on the sky, and correspondingly mapping speed is not at all dependent on sensitivity requirements, rather it depends solely on the receiver FOV and antenna slew rates. See Sect. 3.3 for more considerations regarding multi-pixel detectors.
@@
Figure 4: Left: AtLAST beam diameter as a function of observing frequency. The vertical dashed lines mark frequencies of 90 GHz and 660 GHz. Right: The resulting diameter of the instantaneous FOV as a function of number of detector pixels and observing frequency (see color legend). The resulting ratio of the FOV area and the whole disk of the Sun is overlaid as red contours. The dashed lines mark selected frequencies and detector numbers for reference. The pixels are assumed to have the size of the corresponding beam and are packed into a circle with no overlap. See Sect. 3.3 for details.
@@
A selected target, e.g. a sunspot, would be followed in time for one or several hours (compensating for motion on the sky plus solar rotation) and observed with preferably high cadence. For comparison, ALMA currently observes the Sun at 1 s cadence. This mode, which is equivalent to observations at visible wavelengths at limited field of view, is suitable for a large range of science cases that require a combination of high spatial and temporal resolution.
@@
As a compromise between full-disk scans (mode 1) and single-target tracking (mode 2), AtLAST could perform scans of extended regions (or mosaics). The size of the covered region on the Sun would depend on the instantaneous field of view of the multi-pixel detector (see Sect. 3.3) and the maximum acceptable temporal cadence. Such scans could be repeated continously for some time, e.g. one hour, providing good temporal cadence of a large field of view. This mode could be useful for observations of large-scale features such as filaments or entire Active Regions. Such observations would facilitate the detection of flares and filament eruptions and at the same time capture their evolution at adequate temporal resolution.

### v3

@@ Introduction
Measuring the solar radius at subterahertz frequencies (submm/mm) allows one to probe the solar atmosphere, since these measurements show the height above the photosphere at which most of the emission at a determined observation frequency is generated (Menezes & Valio 2017). Changes in the solar radius show it can be modulated with the 11-year activity cycle (mid-term variations) as well as longer periods (long-term variations). The equatorial radius time series was found to be positively correlated to the solar cycle, since the equatorial regions are more affected by the increase of active regions during solar maxima, making the solar atmosphere warmer in these regions; on the other hand, the anticorrelation between polar radius time series and the solar activity proxies could be explained by a possible increase of polar limb brightening during solar minima (Menezes et al. 2021). In the context of the study on the subterahertz solar atmosphere, spicules can affect the measurement of the solar radius and cause the solar limb to appear more diffuse, which can lead to an overestimation of the solar radius. In Menezes et al. (2022), the discrepancies between measured limb brightening values and model predictions highlight the need for further studies to improve our understanding of the solar atmosphere. Therefore, it is important to take into account the presence of spicules when measuring the solar radius and limb brightening at different frequencies. Currently, however, there are no full-disk observations of the Sun in the mm range apart from occasional ALMA single-dish Total Power (TP) maps a few times a year.
@@
Please note that the Sun’s rotation depends on latitude with a (sidereal) rotation period of down to  24 days at the equator and up to 35 days at the poles (see, e.g., Snodgrass 1983; Balthasar et al. 1986; Schou et al. 1998). Any long-lived feature like an Active Region would thus move across the visible disk of the Sun in less than two weeks. However, at the resolution anticipated for AtLAST (see Figure 3), features like Active Regions and sunspots (Solanki 2003) therein would evolve significantly on a daily timescale. The evolution of Active Regions and the change of the Sun’s magnetic nature across a solar cycle would therefore be addressed best via full-disk mapping of the Sun at multiple frequency bands, once per day. Already an initial 1-month campaign covering a full solar rotation would allow for addressing the fundamental unsolved problems mentioned above including the nature of coronal and chromospheric heating problems, whereas much longer sequences are required for addressing the solar activity cycle. Such sequences would also be relevant for producing a larger sample of flare observations, where the individual observations should be time series over an hour or more as daily maps would not capture the temporal evolution of individual flares (see Sect. 2.2). In addition, these sequences would provide vital complementary input for predictions of the solar activity cycle, which yield important societal impact but, despite much progress, remain challenging (e.g., Hathaway et al. 1999; Dikpati et al. 2006; Cameron & Schüssler 2007; Choudhuri et al. 2007; Pesnell 2012; Balogh et al. 2014).
@@
New insights regarding the Sun’s activity variations and thus the long-term evolution of our host star could be transferred to other stars as demonstrated by stellar radio/mm observations (e.g., Bastian et al. 2000; Selhorst et al. 2013, 2020; Pope et al. 2019). This Sun-as-a-star approach has recently received increasing attention in the context of next-generation exoplanet observations for which it is crucial to separate observable exoplanet signatures from the host star’s “background radiation” (see Rackham et al. 2023, and references therein).
@@
The Sun, which can be observed spatially resolved, serves as a fundamental reference case for studying the dynamic nature of stellar atmospheres and solar/stellar activity including the occurrence of flares (e.g., Dumusque et al. 2015; Strassmeier et al. 2018; Otsu et al. 2022; Zhao et al. 2023; Ma et al. 2024; Pietrow et al. 2024, see also Section 2.2). In addition to serendipitous detection as a by-product of other observations and surveys, dedicated stellar observations are essential for understanding the structure and activity of stellar atmospheres in general. The sample by Mohan et al. (2021) demonstrates the potential of stellar mm observations and, at the same time, illustrates the scarcity of suitable observations. AtLAST could make substantial contributions in this regard, either through targeted observations or also through coincidental (or even archival) detections. In particular, observations of stellar flares in the (sub-)mm range are extremely rare (MacGregor et al. 2018, 2020), and yet are crucial for unveiling the physics of these events including particle acceleration and plasma heating and for evaluating the impact of flares on the habitability of exoplanets (see, e.g., Lammer et al. 2003; Segura et al. 2010; Konings et al. 2022; Bellotti et al. 2023, and references therein). At millimeter wavelengths, the flux density of the few detected stellar flares tends to be 10 - 1000 times higher than the quiescent flux. Assuming a typical quiet chromospheric temperature of $\sim 10^{5}$ K (at 100 GHz) as seen in cool stars (Mohan et al. 2021), the flux density during a flare can vary in the range $\sim$ 0.5 - 50 mJy for a star at a distance of 10 pc. Consequently, stellar flares can be detected even when the stellar disk is spatially unresolved. Observing a broad sample of stellar targets with varying flare intensity distributions therefore complements high-resolution solar flare observations. Translating solar observations into stellar signals, either through dedicated instruments like HARP-N (Cosentino et al. 2012; Dumusque et al. 2015) or POET (Santos 2023), or by processing, i.e. by employing Sun-as-a-star observational techniques, offers significant scientific potential. Such approaches contribute to the study of stellar atmospheres and can enhance our understanding of the physical mechanisms behind solar and stellar flares (see, e.g., Otsu et al. 2022; Ma et al. 2024; Pietrow et al. 2024).
It is important to emphasise that the flux varies on timescales of a few seconds to minutes. Observing the temporal evolution of a flare, which is essential for exploring the physical mechanisms, thus requires time sequences with accordingly high temporal cadence. Such data sets are typically not generated from a serendipitous detection or as part of a survey but require a dedicated sit-and-stare observation of the same region in the sky. Should the required ultra-high temporal resolution not be achievable with full-disk scans of the Sun (see Section 3), high-cadence flare observations of even a limited FOV could be used to transform the data into a Sun-as-a-star observation using a method like, e.g., the one by Pietrow & Pastor Yabar (2023).
@@
The full potential of flare observations at millimeter wavelengths is only unlocked through coordinated, strictly co-simultaneous observations with other telescopes covering as large a wavelength range across the electromagnetic spectrum as possible. Such coordinated campaigns, which have become common in multi-messenger astronomy and are basically standard procedure for observations of the Sun, demand for PI-driven operations. The scarcity of such coordinated stellar observing campaigns is among the main reasons why there has been only a relatively small number of observations that are indicative of a stellar coronal mass ejection (CME) – an essential component of exo-space weather (Houdebine et al. 1990, 1993; Gunn et al. 1994; Guenther & Emerson 1997; Fuhrmeister & Schmitt 2004; Vida et al. 2016; Argiroffi et al. 2019; Veronig et al. 2021; Namekata et al. 2021). The success of such campaigns would be greatly enhanced by AtLAST’s large FOV and high angular resolution, thus being able to observe multiple stellar targets at the same time while separating the flare source(s) from other components such as close-by active companion stars or disks, making AtLAST ideal for time-resolved coordinated observations of stellar flares. See also the time-domain science cases for AtLAST (Orlowski-Scherer & et al. 2024).
@@
Figure 3: The ALMA sunspot mosaic (left) that was taken during the ALMA SV campaign in on December 18, 2015 in Band 6 230 GHz at original interferometric resolution (left) and after convolution with a beam corresponding to a single ALMA 12-m (Total Power) antenna (middle) and after convolution with the expected AtLAST beam (right). For reference, the corresponding beam is plotted in the upper left corner of each panel.
@@
There are two major constraints for solar observations at millimeter wavelengths: (i) As the Sun is by far the brightest millimeter source in the sky, sensitivity is not an issue and very short integration times and thus even on-the-fly scanning is possible. (ii)  The Sun evolves on short time scales down to seconds and below depending on the scientific target, which prohibits long integration/scan times. Imaging of solar data thus needs to be done as snapshot imaging or over short time windows that should not exceed 1 min for the expected resolution (on the Sun, temporal and spatial scales tend to correlate). As will be detailed in this section, the combination of high angular resolution, high temporal cadence (or at least short integration time), wide frequency coverage, and large dynamic range (i.e., a high ratio between the maximum and minimum measurable brightness temperatures) are crucial for meaningful solar observations with AtLAST (Mroczkowski et al. 2024). In addition, full-Stokes polarimetry would greatly enhance AtLAST diagnostic potential for the Sun.
@@
High spatial resolution is of utmost importance for solar observations as much of the relevant dynamics occurs on small spatial scales. Achieving sufficiently high resolution is challenging at millimeter wavelengths, in particular for single-dish telescopes such as AtLAST. While AtLAST’s aperture sets a limit, the achievable resolution is still in the relevant range and superior to any previous single-dish observation of the Sun. With an aperture of 50 m, the achievable angular resolution would be 6.5” at a wavelength of 1.3 mm ($\sim$230 GHz, ALMA Band 6) and down to just 1.5” at 0.3 mm ($\sim$950 GHz, ALMA Band 10) (see Fig.  4).
@@
techniques (with a single beam) are used for creating maps of the whole solar disk or limited regions by, e.g, ALMA (White et al. 2017), and could be implemented for AtLAST, too (Kirkaune et al. in prep.). The difference to mosaics is that the telescope is not consequently pointing at a set of locations on the Sun, possibly staying at each location for a few seconds as done for ALMA interferometric mosaics, but that telescope keeps observing while moving across the Sun to acquire a time stream that can then be converted into a map. This scanning motion could be accomplished by moving the primary or secondary reflector. Scanning the whole disk with an ALMA TP antenna in a double-circle pattern (Phillips et al. 2015) takes on the order of 10 min. See Fig. 1 for examples. The roughly 4 times larger aperture of AtLAST compared to an ALMA 12-m antenna results in an accordingly much smaller ($\sim$6 %) beam area. Scanning the whole disk of the Sun would take considerably longer (on the order of ($(\frac{50\,\mathrm{m}}{12\,\mathrm{m}})^{2}\approx 17$ times, i.e. $\sim$3 hours), which would be sufficient for synoptic studies with just one scan per day but prohibitive for science cases that rely on high temporal cadence. In the latter case, scans would then need to be limited to smaller region, although a compromise between size of the mapped region and overall cadence can be made. In either case, for a time series, the scanning could be continuous.
@@
would significantly enhance the diagnostic possibilities for solar observations with AtLAST. Even for a modest setup with a detector with 1000 pixels in a concentric setup with no overlap, spaced by one beam width, the resulting instantaneously covered region on the Sun (hereafter referred to as the instantaneous field of view, FOV) would cover a region with a radius of about 36 beams. While a single AtLAST beam would be a factor $50/12$ smaller than the beam of a 12 m ALMA antenna, using 1000 pixels would effectively make the instantaneous FOV $36\times\frac{12}{50}\approx 9$ times larger. As shown in Fig. 4, the diameter of the FOV would range from $\sim 45$” at 950 GHz (ALMA Band 10) to $\sim 450$” at 100 GHz (ALMA Band 3). Such a FOV would cover a complete sunspot even at the highest frequencies and whole Active Regions at the lower frequencies. The angular resolution of the final imaging products could be increased by means of a rapid small scan pattern (on scales comparable to the beam size and below) that compensates for the above-described pixel spacing. The diameter of the instantaneous FOV would be halved again for a denser pixel setup that allows for Nyquist sampling, which then would not require the mentioned scan pattern. Please see Sect. 3.3 for a detailed discussion of the detector requirements for solar observations.
@@
might offer means to further increase AtLAST angular resolution. However, as the formation height range of the mm continuum (see also Sect. 2.1) depends on the frequency, only a moderate frequency range should be considered. Depending on the receiver design, multi-frequency synthesis could be an option that could be applied flexibly on AtLAST data sets.
@@
The Sun is by far the brightest millimeter source in the sky with typical brightness temperatures of several thousand Kelvin. Consequently, sensitivity is not a concern and the integration time can be very low, allowing for ultra-high cadence, which is of essence for a range of science cases. The integration time can be much below 1 s depending on possible attenuation strategies (e.g., solar filter or detuning). Important for solar observations, however, is a large dynamic range. ALMA observations in Band 3 and 6 show a wide range of brightness temperatures from as low as 3000 K to above 13 000 K (see, e.g., the data sets in the Solar ALMA Science Archive, Henriques et al. 2022). In general, the mean brightness temperature of the Sun depends on the frequency because the atmospheric height range of the mapped layer depends on the frequency and thus on the thermal stratification of the solar atmosphere. In principle, the highest ALMA receiver bands map the transition between the upper photosphere and lower chromosphere (i.e., the classical “temperature minimum” region) with comparatively low temperatures, whereas the lowest ALMA bands (i.e., at higher frequencies) map radiation emerging from the on average much hotter upper chromosphere close to the transition region with its steep jump in temperature. See also Hofmann et al. (2022) for a discussion of brightness temperatures obtained with ALMA in comparison to other (less temperature-sensitive) diagnostics and (Loukitcheva et al. 2019) for an observation of a particularly cool region. In addition to this already wide brightness temperature range, solar flares as a key science case for solar observations with AtLAST (Section 2.2) define the upper end of extreme temperatures that occur on the Sun. Depending on the strength (i.e. overall energy release) of a flare, plasma temperatures exceed $10^{6}$ K or even $10^{7}$ K. A flare thus emits orders of magnitude more radiation than the average (quiescent) Sun as discussed above. It should be noted that the corresponding peak brightness temperature also depends on the observing frequency and the beam size, which is typically too large to resolve the source. From observations of a (very strong) X-class flare with the 1.5m-aperture Solar Submillimeter Telescope (SST, Kaufmann et al. 2008) at 212 GHz an excess brightness temperature of $1.75\times 10^{6}$ K for an assumed source size of 25”, which would remain unresolved (Giménez de Castro et al. 2018). In comparison, AtLAST would have an angular resolution of $\sim$7” at 212 GHz and would thus be able to resolve the source and thus brightness temperatures in excess of $10^{6}$ K. At shorter wavelengths in the sub-mm range, we rely on simulations to estimate the expected brightness of flares (see Figure 2).
@@
Magnetic fields play an essential role in the Sun’s structure, dynamics, and activity. Their impact is clearly visible in Active Regions most notably in the form of sunspots and as large-scale prominences but also the Sun’s Quiet Regions harbour magnetic fields in various configurations and strengths. The latter becomes immediately obvious from high-resolution H observations (see, e.g., Rutten et al. 2008; Rouppe van der Voort et al. 2023). High-resolution magnetograms obtained from the spectral lines in the photosphere (i.e., the Sun’s surface layer) have become a staple in contemporary solar physics relevant to a large range of scientific topics but often also just for context. Reliably measuring the magnetic field in the solar chromosphere above, i.e. the layer that would be observed with AtLAST, is still a current technological challenge, again with much diagnostic potential (see, e.g., Pietrow et al. 2020; Vissers et al. 2021; Morosin et al. 2022). The measured polarization state provides a measure for the longitudinal component of the magnetic field vector in the same layer as the continuum radiation. In the same way, a scan through wavelength (and thus through formation height) can be used to reconstruct the three-dimensional magnetic field structure in the solar chromosphere. Measuring the magnetic field in this layer is in itself a hot topic with potentially ground-breaking results.
Early observations with ALMA at a frequency of 100 GHz, which were obtained as science verification data in Band 3 (ADS/JAO.ALMA#2011.0.00011.E), indicate that the circular polarization degree is at least 1% . This value is considered a minimum, as the observations were carried out with the 12m-Array only (Shimojo et al. 2024). First regular ALMA observations of the Sun with full polarisation in Band 3 were carried out in 2024 (Cycle 10). Important lessons are expected from these observations, which would have implications for full polarisation observations with AtLAST. In general, circular polarization levels of up to 5% are expected for the Sun at mm wavelengths.
@@
In order to address the aforementioned key science cases (see Sect. 2), it would be best to combine 1) long-term synoptic observations with a very low daily load on observing time with 2) campaign-based observations. The synoptic observations could consist of one daily full-disk map or preferably a 10-30 min long time sequence of full-disk maps across a wide predefined frequency range. The resulting long-term data set would be of essential value for studying the solar cycle and all related variations imprinted in the Sun’s chromosphere with fundamental implications for stellar activity cycles and their impact on the habitability of exoplanets (White et al. 2023). The campaign-based observations would be more adjusted towards the requirements of individual science cases. The scheduling for observations that require the existence of an Active Region could be planned with just a few days ahead as the co-ordination with other space-borne and ground-based observatories would boost the scientific impact of such AtLAST data. Alternatively, a ToO Target-of-Opportunity (ToO) mode (also possibly confined to a campaign) would ensure observations of flares whenever a suitable Active Region appears.
@@
A field-of-view (FOV) that would cover the whole Sun plus a margin around, which is important to include the Sun’s outer atmosphere, has a diameter of $\sim 2000$”. Scanning such a region with the comparatively small beam size of AtLAST would require long scan times (see Sect.  3.1.1). Therefore we expect that multi-pixel detectors will be needed for solar disk mapping as well as other non-solar large FOV studies. The corresponding time needed for scanning the whole FOV would be reduced basically by the number of simultaneously used beams (pixels) and thus the resulting size of the instantaneously covered FOV (see also Sect. 3.1.1). The required scan time for the whole solar disk would scale accordingly with respect to the single-dish Total Power (TP) scans with ALMA antennas, which take about $\sim 10$ min. Consequently, a full-disk scan at the lower frequencies could potentially be completed in a few minutes or less, aided by the anticipated high slew speed of 3 deg/s. A smaller number of pixels, e.g. $\sim 250$ pixels, would directly affect the covered FOV, resulting in correspondingly longer scan times. Please note that very short integration times are sufficient as the Sun is the brightest mm source on the sky, and correspondingly mapping speed is not at all dependent on sensitivity requirements, rather it depends solely on the receiver FOV and antenna slew rates. See Sect. 3.3 for more considerations regarding multi-pixel detectors.
@@
Figure 4: Left: AtLAST beam diameter as a function of observing frequency. The vertical dashed lines mark frequencies of 90 GHz and 660 GHz, which we here define as a minimum requirement for the frequency range. Right: The resulting diameter of the instantaneous FOV as a function of number of detector pixels and observing frequency (see color legend). The resulting ratio of the FOV area and the whole disk of the Sun is overlaid as red contours. The dashed lines mark selected frequencies and detector numbers for reference. The pixels are assumed to have the size of the corresponding beam and are packed into a circle with no overlap. See Sect. 3.3 for details.
@@
A selected target, e.g. a sunspot, would be followed in time for one or several hours (compensating for motion on the sky plus solar rotation) and observed with preferably high cadence ($\sim$1 s). For comparison, ALMA currently observes the Sun at 1 s cadence, demonstrating the scientific potential of such a high cadence, e.g. for the study of waves (e.g., Patsourakos et al. 2020; Eklund et al. 2021; Guevara Gómez et al. 2021; Jafarzadeh et al. 2021). This mode, which is equivalent to observations at visible wavelengths in a limited field of view, is suitable for a large range of science cases that require a combination of high spatial and temporal resolution.
@@
As a compromise between full-disk scans (mode 1) and single-target tracking (mode 2), AtLAST could perform scans of extended regions (or alternatively mosaics, see Section 3.1.1). The size of the covered region on the Sun would depend on the instantaneous field of view of the multi-pixel detector (see Section 3.3) and the maximum acceptable temporal cadence. Such scans could be repeated continuously for some time, e.g., one hour, providing a good temporal cadence of a large field of view. This mode could be useful for observations of large-scale features such as filaments or entire Active Regions. Such observations would facilitate the detection of flares and filament eruptions and, at the same time, capture their evolution at adequate temporal resolution.